# Primordial black hole formation from self-resonant preheating?

### v1

@@ III.1 T-models
The approximate plateau for $\phi\gg M$ is consistent with a power spectrum of primordial curvature fluctuations at CMB scales compatible with observations. In particular, the observational upper limit on the tensor-to-scalar ratio, $r<0.032$ at 95% CL Tristram:2021tvh , requires $M\lesssim 5\div 10\,M_{\rm Pl}$, with the precise maximum allowed value of $M$ depending on the time elapsed between CMB scales leaving the horizon to the end of inflation.
@@
Figure 2: Comparison of the anharmonic term (blue) and the leading metric term (red) for the T-attractor potential assuming $M=7.75\times 10^{-3}M_{\rm P}$ for the numerical solution. The analytical approximation corresponding to eq. (10) for each of the terms is indicated by a black dashed line.
Approximating the end of inflation by the condition $\epsilon_{V}=1$, we get $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$ for $M\gtrsim M_{\rm P}$, and $\phi_{\rm end}\simeq 0.5M\ln(4\sqrt{2}M_{\rm P}/M)$ for $M\ll M_{\rm P}$, which suggests the initial post-inflationary range of interest for the field, $\phi_{\rm end}$, to be roughly given by the interval $[M,M_{\rm P}]$. In Fig. 2 we compare the anharmonic and the dominant metric terms (see eq. (9)): ${V^{\prime\prime}}/{m^{2}}-1$ and $2V^{\prime}\dot{\phi}/(H\,m^{2}M_{\rm P}^{2})$ during the first 3 post-inflationary e-folds, $\Delta N=0$ corresponding to $\epsilon=1$, with field values roughly spanning the range $\phi\in[0.024,1.13]M$, for the choice $M=7.75\times 10^{-3}M_{\rm P}$ (the rationale for the choice $M\ll M_{\rm P}$ will be soon clear). Here we denote with a prime the derivative with respect to $\phi$. The colored lines correspond to the numerical solution, while the dashed ones refer to the analytical approximation valid at late times given by eq. (10). The anharmonic term is much larger than the metric one, and therefore metric preheating is not occurring for this choice of parameters.
@@ III.2 E-models
which, for $n=1$, has the same small-$\phi$ and large $\phi$ limits as in eq. (18) above, and coincides with Starobinsky’s model of inflation Starobinsky:1980te for the choice $M=\sqrt{3/2}M_{\rm Pl}$. Using $\epsilon_{V}=1$ to find the time of the approximate end of inflation, we obtain $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$ for $M\gtrsim M_{\rm P}$ and $\phi_{\rm end}\simeq M\ln(\sqrt{2}M_{\rm P}/M)$ for $M\ll M_{\rm P}$, which again suggests the initial post-inflationary range of interest for the field, $\phi_{\rm end}$, to be roughly given by $[M,M_{\rm P}]$; the same as for T-models. The effective mass squared in the small-field limit is again given by eq. (20). In Fig. 3 we compare the anharmonic and the dominant metric terms, for $M=M_{\rm P}$, which is complementary to the choice made for Fig. 3. We find now that the anharmonic terms are comparable, still marginally larger than the metric ones. Metric-term dominance would require $M\gtrsim M_{\rm P}$, which is only allowed in a small parameter space while remaining phenomenological viable.
Furthermore, even in this case eq. (10) is valid only at late times while it is a poor approximation of the dynamics just after the end of inflation. We thus conclude that the quadratic approximation is nowhere reliable at a quantitative level. The same argument applies also to the $\alpha$-attractor T-models.
Figure 3: Comparison of the anharmonic term (blue) and the leading metric term (red) for the E-attractor potential assuming $M=M_{\rm P}$, close to the popular Starobinsky model, for the numerical solution. The analytical approximation corresponding to eq. (10) is indicated by a black (anharmonic) and gray (metric) dashed line.
@@ III.3 The possible role of anharmonic terms
We have seen in the previous section that current knowledge suggests that a quadratic potential does not lead to PBH formation from preheating. However, the results of this section may open an opportunity: Since in most of the parameter space anharmonic terms are at least as important, and often dominant, with respect to metric ones, maybe one can rescue the idea of forming PBHs via a self-resonant instability, when anharmonicity is accounted for.
@@ IV Non-linear evolution: lattice results
Based on the previous considerations, the most promising case to consider for PBH formation from self-resonant preheating is the one characterised by $M\ll M_{\rm P}$, on which we will now focus.999 Note that the complementary, intermediate regime $0.01\lesssim M\lesssim 0.1$, which experience only moderate resonances, has been recently studied in Eggemeier:2023nyu . In this case, a strong resonance regime ($\Re(\mu_{k})m/H\gtrsim 7$) is responsible for the amplification of the perturbations. Furthermore, we have shown that in this case the dynamics leading to parametric amplification is dominantly driven by the anharmonic terms in the potential, rather than the metric perturbations.
@@
Metric perturbations are neglected in these lattice simulations. In the linear regime, this is justified for the parameter space of interest, as shown in sec. II. As extensively argued in Ref. Lozanov:2019ylm , this is also the case at later stages of the evolution since the Newtonian potential $\Phi$ remains small. In particular, the gravitational potential on the surface of oscillons and transients is estimated to be $|\Phi|\lesssim 10^{-3},$ thus far from the values $|\Phi|\sim\mathcal{O}(1)$ required to form a BH. Hence, for $M\ll M_{\rm Pl}$ the parametric growth of the inflaton perturbations during preheating leads to the formation of weakly gravitating metastable objects, which are not expected to collapse into PBHs, rather to eventually dissipate. This suggests that the production of PBHs from preheating is far from generic in phenomenologically viable, well-motivated, and simple inflationary scenarios, such as the one we have considered.
As briefly discussed in Sec. II.1, PBH formation is at least in principle possible when $M\sim M_{\rm Pl}$, as suggested by numerical relativity simulations in Muia:2019coe ; Nazari:2020fmk ; Kou:2019bbc (see also Padilla:2021zgm ; Hidalgo:2022yed for related analytical work). However, the question of whether the needed conditions for PBH formation are generically found in self-consistent inflationary models is rarely addressed. Instead, parametrically set initial conditions for the inflaton overdensity are typically adopted. For instance, quoting from Ref. Muia:2019coe :
@@
In Fig. 5 we do perform such a test with a lattice calculation, similar to the case shown in Fig. 4, but adopting $M=0.2M_{\rm P}$ to mimic the parameter choice adopted in Muia:2019coe . Note that for this choice, the metric backreaction is still sub-leading, and the lattice results should still be at least qualitatively reliable. Our results show that the enhancement of the power-spectrum, while present, is many orders of magnitude weaker compared to what would be needed for the simulations of ref. Muia:2019coe leading to PBH formation to be considered very representative.
Although these tests are not exhaustive, they all confirm our expectation that, in agreement with intuition, very specific conditions on the inflaton potential are actually needed to trigger PBH formation in the post-inflationary era.
@@ V Conclusions
As a preliminary remark, it is worth stressing that the fluctuation enhancement one can obtain is at best [^10] This has to do with two facts: i) the average equation of state is slightly larger than zero, so that the growth of fluctuations is slightly less efficient than in matter domination, since $\delta\propto a^{1-3w}$ and $\langle w\rangle>0$ (see Appendix A); ii) generalised Floquet theory shows that only some wavelength bands, among all the ones allowed by the Jeans criterion, are actually allowed to grow, see the discussion around eqs. (15),(16). equivalent to the standard Newtonian growth of perturbations in a matter-dominated expansion phase. A scrutiny of the literature shows that forming PBHs in a matter-driven expansion phase is much more challenging and definitely not generic. We attribute this discrepancy to a too simplistic treatment of the non-linear growth of perturbations, that neglects non-linear feedback (and have been treated in a too idealised setting), and/or to an unjustified application of horizon-scale (energy density contrast) criteria for collapse into PBHs to sub-horizon fluctuations.
@@
Furthermore, we have argued that a treatment of metric preheating independent of the specifics of inflaton potential, i.e. limited to the quadratic approximation, can not be generically used for reliable quantitative studies. In fact, in realistic models anharmonic terms in the potential can be sizeable and dominate the dynamics describing the evolution of the density perturbations. For the sake of concreteness, we have demonstrated this result in the context of the $\alpha$-attractor T-models and E-models. In general, we could not identify a situation where the quadratic approximation and the metric preheating dominance are simultaneously verified.
@@
Based on this insight, we have investigated with lattice calculations if the anharmonicity can be used to rescue the original idea, i.e. to resort to inflaton self-interactions to boost the formation of non-linear structures, thus easing PBH formation. We find that this is not the case, since the parameter space where the formation of non-linear structures is significantly enhanced (corresponding to the quadratic to flat potential transition scale $M$ being much smaller than $M_{\rm P}$) corresponds to the one where these objects are metastable oscillons, rather than PBHs. We have recovered results presented in the context of early universe oscillon physics and rarely appreciated in the context of PBH literature.
@@
Based on a few GR numerical studies Muia:2019coe ; Nazari:2020fmk ; Kou:2019bbc ; deJong:2021bbo , the least challenging parameter space to form PBHs by self-resonance in the post-inflationary phase seems to correspond to the range $M\lesssim M_{\rm P}$. We have raised awareness on the fact that even in this case, it may be difficult to obtain the needed large inhomogeneous initial conditions from the mere quantum fluctuations induced from generic and phenomenologically viable inflationary potentials subject to parametric amplification. A further exploration of self-consistent scenarios fulfilling these conditions certainly deserves further study, but it is already clear that PBH formation is not generic even in this case.
@@
Despite the largely negative nature of the results we obtained, we would like to stress that the post-inflationary scalar field dynamics remains a rich and interesting topic, which has been explored and still deserves further exploration concerning a number of phenomena, such as primordial gravitational waves, magnetic fields, non-gaussianities, baryogenesis, and dark matter, beyond the topic of PBH production Amin:2014eta . Even limiting oneself to PBHs, besides further numerical GR studies, another possibility to form sub-horizon scale PBH relies on the interaction of oscillons/solitons both via gravity and new interactions, as reviewed in Flores:2024eyy . Existing studies on gravitational scattering of oscillons/solitons after preheating show some interesting dynamical clustering, but do not provide any indication of PBH formation, suggesting that if PBH formation happens, it must be a rare process Amin:2019ums . We would then argue that our conclusion that PBH formation is not generic should be read as an opportunity, in case of PBH detection, to shed light on possible peculiar dynamics in the early Universe.
@@ Appendix A Post-inflationary matter-like expansion: A closer inspection
We analyse the equation of state and the sound speed associated to the homogeneous inflaton field oscillating around the minimum of a quadratic potential after inflation. Considering the asymptotic behaviour in eq. (10), the energy density $\rho(t)$ and pressure density $P(t)$ are:
@@
In practice, these simple analytical averages do not capture the leading corrections to the approximation $\rho=\bar{\rho}$ and $P=0$. In fact one should take into account that i) the scale factor evolves approximately as $a(t)\sim t^{2/3}$, ii) the solution in eq. (10) is approximate. Both effects are of the order ${\cal O}(H/m)$, i.e. larger than the correction obtained above Cembranos:2015oya .

### v3

@@ III.1 T-models
[…]
@@
Figure 2: Comparison of the anharmonic term (blue) and the leading metric term (red) for the T-attractor potential assuming $M=7.75\times 10^{-3}M_{\rm P}$ for the numerical solution. The analytical approximation corresponding to eq. (10) for each of the terms is indicated by a black dashed line, where we numerically solved for the scale factor.
Approximating the end of inflation by the condition $\epsilon_{V}=1$, we get $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$ for $M\gtrsim M_{\rm P}$, and $\phi_{\rm end}\simeq 0.5M\ln(4\sqrt{2}M_{\rm P}/M)$ for $M\ll M_{\rm P}$, which suggests the initial post-inflationary range of interest for the field, $\phi_{\rm end}$, to be roughly given by the interval $[M,M_{\rm P}]$. In Fig. 2 we compare the anharmonic and the dominant metric terms (see eq. (9)): ${V^{\prime\prime}}/{m^{2}}-1$ and $2V^{\prime}\dot{\phi}/(H\,m^{2}M_{\rm P}^{2})$ during the first 3 post-inflationary e-folds, $\Delta N=0$ corresponding to $\epsilon=1$, with field values roughly spanning the range $\phi\in[0.024,1.13]M$, for the choice $M=7.75\times 10^{-3}M_{\rm P}$ (the rationale for the choice $M\ll M_{\rm P}$ will be soon clear). Here we denote with a prime the derivative with respect to $\phi$. The colored lines correspond to the numerical solution, while the dashed ones refer to the analytical approximation valid at late times given by eq. (10). The anharmonic term is much larger than the metric one, and therefore metric preheating is not occurring for this choice of parameters. It is important to stress that the evolution depicted in Fig. 2 cannot be meaningfully extrapolated to many e-folds. Both the exponential amplification of the inflaton perturbations and the subsequent backreaction (as we will see) take place within one e-fold after inflation (see Fig. 4 later on), at times where the assumption that the inflaton is undergoing damped oscillations in a quadratic potential is not justified.
@@ III.2 E-models
which, for $n=1$, has the same small-$\phi$ and large $\phi$ limits as in eq. (18) above, and coincides with Starobinsky’s model of inflation Starobinsky:1980te for the choice $M=\sqrt{3/2}M_{\rm P}$. Using $\epsilon_{V}=1$ to find the time of the approximate end of inflation, we obtain $\phi_{\rm end}\simeq\sqrt{2}M_{\rm P}$ for $M\gtrsim M_{\rm P}$ and $\phi_{\rm end}\simeq M\ln(\sqrt{2}M_{\rm P}/M)$ for $M\ll M_{\rm P}$, which again suggests the initial post-inflationary range of interest for the field, $\phi_{\rm end}$, to be roughly given by $[M,M_{\rm P}]$; the same as for T-models. The effective mass squared in the small-field limit is again given by eq. (20). In Fig. 3 we compare the anharmonic and the dominant metric terms, for $M=M_{\rm P}$, which is complementary to the choice made for Fig. 3. We find now that the anharmonic terms are comparable, still marginally larger than the metric ones. Metric-term dominance would require $M\gtrsim M_{\rm P}$, which is only allowed in a small parameter space while remaining phenomenological viable.
Furthermore, even in this case eq. (10) is valid only at late times while it is a poor approximation of the dynamics just after the end of inflation, where however the instability is already operational. We thus conclude that the quadratic approximation is nowhere reliable at a quantitative level. The same argument applies also to the $\alpha$-attractor T-models.
Figure 3: Comparison of the anharmonic term (blue) and the leading metric term (red) for the E-attractor potential assuming $M=M_{\rm P}$, close to the popular Starobinsky model, for the numerical solution. The analytical approximation corresponding to eq. (10) is indicated by a black (anharmonic) and gray (metric) dashed line, where we numerically solved for the scale factor.
@@ III.3 The possible role of anharmonic terms
We have seen in the previous section that current knowledge suggests that a quadratic potential (recovered in our potentials when $M\gg M_{\rm P}$, being careful at the limited regime where it is viable) does not lead to PBH formation from metric-preheating. However, the results of this section may open an opportunity: Since in most of the parameter space anharmonic terms are at least as important (when $M\sim M_{\rm P}$), and often dominant (when $M<M_{\rm P}$), with respect to metric ones, maybe one can rescue the idea of forming PBHs via a self-resonant instability, when anharmonicity is accounted for.
@@ IV Non-linear evolution: lattice results
Based on the previous considerations, the most promising case to consider for PBH formation from self-resonant preheating is the one characterised by $M\ll M_{\rm P}$, on which we will now focus.999 Note that the complementary, intermediate regime $0.01\lesssim M/M_{P}\lesssim 0.1$,which experience only moderate resonances, has been recently studied in Eggemeier:2023nyu . Notice that for sufficiently smaller values of $M$ tachyonic preheating may become more dominant than metric preheating, see Tomberg:2021bll . In this case, a strong resonance regime ($\Re(\mu_{k})m/H\gtrsim 7$) is responsible for the amplification of the perturbations. Furthermore, we have shown that in this case the dynamics leading to parametric amplification is dominantly driven by the anharmonic terms in the potential, rather than the metric perturbations.
@@
Metric perturbations are neglected in these lattice simulations. In the linear regime, this is justified for the parameter space of interest, as shown in sec. II. As extensively argued in Ref. Lozanov:2019ylm , this is also the case at later stages of the evolution since the Newtonian potential $\Phi$ remains small. This can be understood making use of its relation to density fluctuations for modes that are deeply sub-Hubble: $\Phi\sim(aH/k)^{2}\delta\ll\delta$, where $\delta$ is the density perturbation. Importantly, since we are dealing with modes for which $k\gg aH$, this relation is gauge independent. So is the variable $\delta\phi$ describing the fluctuations of the inflaton field, which in this regime is just another way of expressing the Mukhanov-Sasaki variable $v$. In particular, the gravitational potential on the surface of oscillons and transients is estimated to be $|\Phi|\lesssim 10^{-3},$ thus far from the values $|\Phi|\sim\mathcal{O}(1)$ required to form a BH. Hence, for $M\ll M_{\rm P}$ the parametric growth of the inflaton perturbations during preheating is important, and, as we already discussed in Sec. III, it corresponds to the regime where anharmonic terms dominate over metric contributions, justifying resorting to CL, where the latter are neglected. In this regime, parametric instability leads to the formation of weakly gravitating metastable objects, which are not expected to collapse into PBHs, rather to eventually dissipate. This suggests that the production of PBHs from preheating is far from generic in phenomenologically viable, well-motivated, and simple inflationary scenarios, such as the one we have considered.
PBH formation is at least in principle possible when $M\sim M_{\rm P}$, as suggested by numerical relativity simulations in Muia:2019coe ; Nazari:2020fmk ; Kou:2019bbc (see also Padilla:2021zgm ; Hidalgo:2022yed for related analytical work). However, the question of whether the needed conditions for PBH formation are generically found in self-consistent inflationary models is rarely addressed. Instead, parametrically set initial conditions for the inflaton overdensity are typically adopted. For instance, quoting from Ref. Muia:2019coe :
@@
In Fig. 5 we do perform such a test with a lattice calculation, similar to the case shown in Fig. 4, but adopting $M=0.2M_{\rm P}$ to mimic the parameter choice adopted in Muia:2019coe . Note that for this choice, the metric backreaction is still sub-leading, and the lattice results should still be at least qualitatively reliable. Our results show that the enhancement of the power-spectrum, while present, is $\sim 8$ orders of magnitude weaker compared to what would be needed for the simulations of ref. Muia:2019coe leading to PBH formation to be considered very representative. Indeed $\mathcal{O}(1)$ values of the metric potential are required for collapse into a black hole.
[…]
At the current state of knowledge, all the indications confirm our expectation that, in agreement with intuition, very specific conditions on the inflaton potential are actually needed to trigger PBH formation in the post-inflationary era: The regime where parametric self-resonance is prominent is in tension with the requirement of forming PBHs, at least in single field models.
@@ V Conclusions
As a preliminary remark, it is worth stressing that the fluctuation enhancement one can obtain is at best [^10] This has to do with two facts: i) the average equation of state is slightly larger than zero, so that the growth of fluctuations is slightly less efficient than in matter domination, since $\delta\propto a^{1-3w}$ and $\langle w\rangle>0$ (see Appendix A); ii) generalised Floquet theory shows that only some wavelength bands, among all the ones allowed by the Jeans criterion, are actually allowed to grow, see the discussion around eqs. (15),(16). equivalent to the standard Newtonian growth of perturbations in a matter-dominated expansion phase. A scrutiny of the literature shows that forming PBHs in a matter-driven expansion phase is much more challenging and definitely not generic. We attribute this discrepancy with the original publications on the topic to a too simplistic treatment of the non-linear growth of perturbations, that neglects non-linear feedback (and have been treated in a too idealised setting); in some other articles, much more optimistic conclusions can be traced back to an unjustified application of horizon-scale (energy density contrast) criteria for collapse into PBHs to sub-horizon fluctuations.
@@
Furthermore, we have argued that a treatment of metric preheating independent of the specifics of inflaton potential, i.e. limited to the quadratic approximation, can not be generically used for reliable quantitative studies. In fact, in realistic models anharmonic terms in the potential can be sizeable and dominate the dynamics describing the evolution of the density perturbations. For the sake of concreteness, we have demonstrated this result in the context of the $\alpha$-attractor T-models and E-models. In general, we could not identify a situation where the quadratic approximation and the metric preheating dominance are simultaneously verified to quantitatively rely on the approximate treatment common in the literature.
@@
Based on this insight, we have investigated with lattice calculations if the anharmonicity can be used to rescue the original idea, i.e. to resort to inflaton self-interactions to boost the formation of non-linear structures, thus easing PBH formation. We find that this is not the case, since the parameter space where the formation of non-linear structures is significantly enhanced (corresponding to the quadratic to flat potential transition scale $M$ being significantly smaller than $M_{\rm P}$, and where GR effects are unimportant) corresponds to the one where these objects are metastable oscillons, rather than PBHs. We have recovered results presented in the context of early universe oscillon physics and rarely appreciated in the context of PBH literature.
@@
Based on a few GR numerical studies Muia:2019coe ; Nazari:2020fmk ; Kou:2019bbc ; deJong:2021bbo , the least challenging parameter space to form PBHs by self-resonance in the post-inflationary phase seems to correspond to the range $M\lesssim M_{\rm P}$. We have raised awareness on the fact that even in this case, it may be difficult to obtain the needed large inhomogeneous initial conditions from the mere quantum fluctuations induced from generic and phenomenologically viable inflationary potentials subject to parametric amplification, confirming the results recently presented in Aurrekoetxea:2023jwd . A further exploration of self-consistent scenarios fulfilling these conditions certainly deserves further study, but it is already clear that PBH formation is not generic even in this case.
@@
Despite the largely negative nature of the results we obtained, we would like to stress that the post-inflationary scalar field dynamics remains a rich and interesting topic, which has been explored and still deserves further exploration concerning a number of phenomena, such as primordial gravitational waves, magnetic fields, non-gaussianities, baryogenesis, and dark matter, beyond the topic of PBH production Amin:2014eta . Even limiting oneself to PBHs, besides further numerical GR studies, another possibility to form sub-horizon scale PBH relies on the interaction of oscillons/solitons both via gravity and new interactions, as reviewed in Flores:2024eyy . Existing studies on gravitational scattering of oscillons/solitons after preheating show some interesting dynamical clustering, but do not provide any indication of PBH formation, suggesting that if PBH formation happens, it must be a rare process Amin:2019ums . We would then argue that our conclusion that PBH formation is not generic should be read as an opportunity, in case of PBH detection, to shed light on very peculiar dynamics in the early Universe.
@@ Appendix A Post-inflationary matter-like expansion: A closer inspection
Depending on the applications, the approximation of using the perfect fluid, dust-like equation of state for the scalar field density and pressure may prove inadequate, and the actual dynamics may deserve going beyond that approximation. See for instance Martin:2020fgl for an example of the changes associated to the presence of the radiation background the inflaton starts decaying into. Here, we analyse the equation of state and the sound speed associated to the homogeneous inflaton field oscillating around the minimum of a quadratic potential after inflation. Considering the asymptotic behaviour in eq. (10), the energy density $\rho(t)$ and pressure density $P(t)$ are:
@@
In practice, these simple analytical averages do not capture the leading corrections to the approximation $\rho=\bar{\rho}$ and $P=0$. In fact, one should take into account that i) the scale factor evolves approximately as $a(t)\sim t^{2/3}$, ii) the solution in eq. (10) is approximate. Both effects are of the order ${\cal O}(H/m)$, i.e. larger than the correction obtained above Cembranos:2015oya .